# 3DXTalker: Unifying Identity, Lip Sync, Emotion, and Spatial Dynamics in Expressive 3D Talking Avatars

### v2

@@ Abstract
Audio-driven 3D talking avatar generation is increasingly important in virtual communication, digital humans, and interactive media, where avatars must preserve identity, synchronize lip motion with speech, express emotion, and exhibit lifelike spatial dynamics, collectively defining a broader objective of expressivity. However, achieving this remains challenging due to insufficient training data with limited subject identities, narrow audio representations, and restricted explicit controllability. In this paper, we propose 3DXTalker, an expressive 3D talking avatar through data-curated identity modeling, audio-rich representations, and spatial dynamics controllability. 3DXTalker enables scalable identity modeling via 2D-to-3D data curation pipeline and disentangled representations, alleviating data scarcity and improving identity generalization. Then, we introduce frame-wise amplitude and emotional cues beyond standard speech embeddings, ensuring superior lip synchronization and nuanced expression modulation. These cues are unified by a flow-matching-based transformer for coherent facial dynamics. Moreover, 3DXTalker also enables natural head-pose motion generation while supporting stylized control via prompt-based conditioning. Extensive experiments show that 3DXTalker integrates lip synchronization, emotional expression, and head-pose dynamics within a unified framework, achieves superior performance in 3D talking avatar generation.
Figure 1: Overview of our expressive 3D talking avatar generation. Given a single reference image and a driving speech audio, 3DXTalker generates identity-consistent 3D talking avatars with accurate lip synchronization, expressive emotions, and natural head-pose dynamics.
@@ 1 Introduction
Audio-driven 3D talking avatar generations [56, 4, 51, 26] have been extensively applied across various domains. By mapping audio signals (speech or singing) to realistic 3D facial movements, these systems provide a natural way to animate avatars without requiring complex 3D capture hardware. Early approaches [17, 13, 38] largely focused on producing basic lip movements or coarse facial animation, which underperformed against the emerging demand for personalized, natural move and communicative avatars. As applications evolve, the field has shifted toward a more comprehensive goal: achieving expressivity, where avatars could preserve identity, synchronize lips with speech, express emotions, and exhibit lifelike spatial, shown as Figure 1.
Despite notable progress [15, 35, 25, 42], achieving comprehensive expressivity remains an ongoing challenge, primarily due to three key factors: insufficient training data with limited identity diversities, narrow audio representations, and restricted explicit controllability. Existing 3D audio–mesh datasets [12, 49, 18] rely on costly, complex real-world capture, while many modeling paradigms [16, 17, 35] couple identity with motion patterns, resulting in limited identity diversity, emotion, and spatial motions, which hinders the models’ scalability and generalization. Meanwhile, commonly used audio representations mainly capture linguistic content [23, 7] (e.g., Wav2Vec [3]), while under-representing prosodic cues, e.g., amplitude variation and emotional tone, are critical for accurate lip movements and expressive facial emotions. Moreover, explicit controllability is often limited, and most talking heads are generated in a largely static or front-facing configuration and ignore camera motion during rendering [28, 1, 48], constraining the ability of user-defined spatial dynamics. These limitations collectively hinder a model’s ability to capture the full spectrum of speech-driven expressivity for identity, lip sync, emotion, and spatial dynamics.
Generating identity-diverse 3D talking heads is challenging because high-quality 3D data is scarce, while 2D videos provide abundant identities, emotional styles, and natural motion patterns at scale. Facial representation models, such as DECA [19] and EMOCA [14], could lift video frames into FLAME shape–expression–pose parameters [27], effectively bridging 2D video and controllable 3D facial modeling. Building on this paradigm, we construct a 2D-to-3D data-curated identity modeling pipeline. We collect three lab-controlled datasets (GRID [11], RAVDESS [29], MEAD [46]) and three in-the-wild datasets (VoxCeleb2 [10], HDTF [52], CelebV-HQ [53]). A unified filtering pipeline is then applied for duration, language, audio–visual sync, and resolution to ensure data quality and diversity. We build structured frame-referred FLAME parameters using EMOCA, leveraging its disentangled latent space to separate stable identity attributes from temporally varying lip, expression, and pose motions. This yields a dataset with rich identity coverage, lip and emotional variations, and lifelike spatial dynamics, without relying on 3D capture.
@@
Utilizing the above pipeline, we introduce 3DXTalker, an integrated flow-matching framework that operates in the disentangled FLAME parameter space to generate expressive 3D talking-head motion sequences conditioned on a reference image and driving audio. Beyond conventional audio embeddings for lip sync, we incorporate frame-wise amplitude features for coherent mouth aperture and frame-wise emotion features for nuanced expression modulation. This forms audio-rich representations that more faithfully reflect the dynamics of speech. The image-derived identity latent and audio-derived motion cues are jointly modeled via a multi-branch transformer, which produces identity-consistent, emotionally aligned talking-head motion. To further enhance dynamic controllability, we enable scalable adjustment of global expression intensity, and optionally incorporate LLM-driven head-pose modulation. Collectively, these designs make 3DXTalker simultaneously achieve identity consistency, lip-sync accuracy, emotional expressions, and diverse spatial dynamics in a single paradigm.
The main contributions are summarized as follows:
@@
    We construct a scalable 2D-to-3D data-curated identity modeling with disentangled representation, mitigating both the limited dataset and identity diversity problems.
@@
    We introduce frame-wise amplitude and emotion cues beyond standard speech embeddings, to further improve lip synchronization and emotional expressions.
@@
    We integrate these advances into 3DXTalker, a comprehensive system that jointly achieves holistic expressivity in 3D talking avatars, while supporting controllable head-pose generation and emotion diversity.
@@ 2 Related Work
3D Identity Modeling and Avatars. Audio-driven 3D avatar synthesis relies on parametric or non-parametric mesh representations. Parametric models, such as 3D Morphable Model [5], Basel Face Model [21], and FLAME [27], use low-dimensional shape, expression, and pose parameters to represent facial geometry, enabling 2D-to-3D learning methods [19, 14, 20, 55]. In contrast, non-parametric approaches model facial surfaces at the vertex or point-cloud level for greater expressiveness. Building on these representations, recent works [13, 38, 33, 54, 47, 8] animate 3D heads from audio, evolving from deterministic regression mappings to generative models.
@@
Regression Models with Audio. Early methods adopt regression frameworks that deterministically map audio to 3D facial motion. Most extract features from large-scale self-supervised speech models, such as Wav2vec 2.0 [3], HuBERT [23], and WavLM [7], and regress either raw vertex displacements [17, 32, 34, 24] or parametric model latents [25, 31, 43, 16, 50, 35, 15, 39]. FaceFormer [17] predicts vertex trajectories using a transformer decoder, while UniTalker [16] and CodeTalker [50] compress facial motion into low-dimensional latents for efficient prediction.
@@
Generative Models with Audio. To improve the diversity and realism of audio-driven 3D facial animation, recent work has increasingly adopted generative formulations that model the conditional distribution from speech audio to facial motion. VQ-VAE–based methods (e.g., Learn2Listen [31], DEEPTalk [25]) discretize facial motion into latent code sequences and sample them conditioned on audio, while diffusion-based approaches (e.g., FaceDiffuser [41], DiffPoseTalk [42], DiffusionTalker [6], FaceTalk [1]) generate motion by progressively denoising from Gaussian noise, yielding higher variation and realism. Despite these advances in lip synchronization, existing methods typically enhance expressivity along other dimensions. For instance, EMOTE [15] focuses on modeling emotional facial expressions, while DiffPoseTalk explores audio-driven head-pose dynamics as a separate component. We therefore propose an integrated framework that unifies these three components for comprehensively enhancing expressive 3D talking avatar generation.
@@ 3 Method
In the task of 3D audio-driven talking avatar generation, the objective is to synthesize a sequence of 3D avatar states $\{\mathbf{M}_{i}\}_{i=1}^{N}$ (NN total frames) that align with a given audio waveform $\mathbf{A}$, ensuring precise synchronization between facial movements and speech. Beyond lip-sync, modern avatar systems are increasingly expected to achieve expressivity, which encompasses consistent identity preservation, emotion-aware facial expressions, and realistic spatial dynamics that enhance communicative impact. Meanwhile, there remains room to further enhance expressivity by expanding dataset and identity diversity for better generalization, enriching audio representations to capture richer prosodic patterns and emotional cues, and strengthening explicit controllability for more interpretable and precise manipulation of avatar dynamics. To meet the diverse requirements of expressivity, we propose an integrated framework, 3DXtalker, that enhances avatar generation across four dimensions: identity consistency, lip synchronization, emotion expressions, and spatial dynamics controllability, as detailed in the following subsections.
@@
3.1 Data-curated Identity Modeling
EMOCA Modeling Preliminary. To leverage the abundant identities, emotional styles, and dynamic motion patterns in 2D videos, we adopt the EMOCA parametric autoencoder [14], which projects video frames into a controllable 3D facial parameter space. The model encodes 2D images into FLAME parameters [27], including shape $\boldsymbol{\beta}\in\mathbb{R}^{100}$, pose $\boldsymbol{\theta}\in\mathbb{R}^{6}$, expression $\boldsymbol{\psi}\in\mathbb{R}^{50}$ and an additional detail parameter $\boldsymbol{\delta}\in\mathbb{R}^{128}$. Then, they can be decoded into a coarse mesh $\mathbf{M}_{coa}\in\mathbb{R}^{5023\times 3}$ deformed from a FLAME template $\mathbf{\bar{V}}$ with $\boldsymbol{\beta}$, $\boldsymbol{\psi}$ and $\boldsymbol{\theta}$, which is refined by detail decoder $\mathcal{D}_{det}$} to form a detailed mesh $\mathbf{M}_{det}\in\mathbb{R}^{59315\times 3}$ through facial displacement map, formally defined as:
@@ 3.1 Data-curated Identity Modeling
where $\mathbf{V}_{int}$ is an intermediate mesh, $\text{B}_{S}$ and $\text{B}_{E}$ are shaping and expressing functions, LBS denotes the linear blend skinning algorithm with joint regressors $\text{J}_{P}$ and skinning weights $\mathcal{W}$, $\mathbf{\bar{U}}$ denotes template UV map, and $\text{F}_{det}$ applies the predicted displacement map to the coarse mesh $\mathbf{M}_{coa}$. Notably, the pose parameter $\boldsymbol{\theta}$ can be decomposed into head pose and jaw pose, allowing head motion to be modeled and controlled through simple linear variations. This parameterization provides a disentangled and controllable 3D facial representation, which serves as the foundation for our subsequent data-curated identity modeling pipeline. Further details of this process are provided in Appendix A.
Dataset Curation Pipeline. We construct a diverse data corpus by combining three lab-controlled datasets (GRID [11], RAVDESS [29], MEAD [46]) with three in-the-wild datasets (VoxCeleb2 [10], HDTF [52], CelebV-HQ [53]). The lab-controlled datasets provide high-quality recordings with articulated facial movements and emotional expressions, while the in-the-wild datasets introduce a broad range of identities, speaking styles, and natural spatial dynamics. To ensure data reliability across sources, we apply a unified filtering pipeline comprising: (1) duration thresholding to stitch video clips exceeding 10 seconds, (2) signal-to-noise ratio filtering to suppress noisy or corrupted speech, (3) language filtering to maintain linguistic consistency, (4) audio–visual synchronization verification to eliminate misaligned segments, and (5) spatial resolution normalization to $512\times 512$. (Further details provided in Appendix B.1.)
Each video frame is then lifted into the FLAME parameter space via EMOCA [14], yielding a structured representation that consists of identity shape $\boldsymbol{\beta}$ and detail $\boldsymbol{\delta}$, as well as frame-varying expression $\boldsymbol{\psi}$ and head pose $\boldsymbol{\theta}$. To construct a temporally consistent representation that disentangles identity from motion, we take the first frame as the reference and express all parameters in a differential form:
@@
3.2 Unified Audio-rich Framework
Overview. The data-curated identity modeling pipeline obtains disentangled identity and motion representations from 2D videos using EMOCA. This allows a single reference image to serve as a reliable identity anchor during generation, in contrast to prior 3D audio-driven methods that relied on a fixed identity template. Additionally, the audio side remains a limiting factor. Speech signals inherently contain multiple layers of information: (1) linguistic content (word, phoneme), (2) articulatory dynamics reflected in amplitude and rhythm that drive jaw motion and mouth aperture, and (3) emotional prosody conveyed through intonation, energy contours, and vocal timbre. Common audio embeddings focus on linguistic content [3, 23, 7], but overlook prosodic cues, leading to correct word synchronization but weak mouth aperture and flat emotion expressions.
To address this gap, we introduce 3DXTalker, an integrated flow-matching framework that operates in the disentangled parameter space to generate expressive 3D talking-head sequences conditioned on a reference image $\mathbf{I}_{0}$ and driving audio $\mathbf{A}$. Beyond conventional audio embeddings, we incorporate frame-wise amplitude features for coherent mouth aperture and frame-wise emotion features for nuanced expression modulation, forming audio-rich representations that more precisely reflect the dynamics of speech. The identity latent derived from the reference image and the audio cues is then jointly modeled through a multi-branch flow-matching transformer, enabling identity consistency, lip synchronization, and emotional alignment in talking-head motion, as depicted in Figure 2. We next outline the inference process of this framework.
Transformer Backbone. Given an input image $\mathbf{I}_{0}$, we first extract the reference parametric representation $\mathbf{X}_{ref}=(\boldsymbol{\beta}_{0},\boldsymbol{\theta}_{0},\boldsymbol{\psi}_{0},\boldsymbol{\delta}_{0})\in\mathbb{R}^{1\times 284}$, which is then combined with $t$ step-dependent noise $\boldsymbol{\varepsilon}_{t}\in\mathbb{R}^{N\times 284}$ (initial random noise $\boldsymbol{\varepsilon}_{0}$) to form the model’s input state $\mathbf{\tilde{X}}_{t}$ after the MLP layers. The input audio $\mathbf{A}$ is processed into linguistic embeddings $\mathbf{A}_{feat}\in\mathbb{R}^{N\times d}$ ($d$ represents dimension) using WavLM [7]. In the Transformer backbone $\mathcal{D}_{flow}$, $\mathbf{\tilde{X}}_{t}$ serves as queries, while the audio embeddings $\mathbf{A}_{feat}$ act as keys and values. Through self-attention and cross-attention, the model produces the intermediate latent representation $\mathbf{H}_{t}$. The overall process can be summarized as follows:

### v3

@@ Abstract
Audio-driven 3D talking avatar generation is increasingly important in a wide range of multimedia applications. In these scenarios, avatars are expected to preserve identity, synchronize lip motion with speech, convey rich emotions, and exhibit flexible head pose, which together define the overall goal of expressivity. However, achieving this goal remains challenging due to insufficient diversity of identities, expressions, and pose patterns in existing datasets, the absence of an integrated system, and limited semantic controllability. In this paper, we propose 3DXTalker, [^1] We will open-source the complete data integration pipeline, processed dataset, and train-inference codebase forfuture research on expressive 3D talking avatars. an integrated framework for expressive 3D talking avatar generation. First, we develop a scalable 2D-to-3D data integration pipeline that bridges large-scale videos and structured 3D facial representations, improving facial diversity. Second, we introduce a flow-matching framework in a disentangled parameter space, combining image-derived identity with audio-rich representations enhanced by frame-wise amplitude and emotion cues, to generate expressive 3D facial motions. Finally, we adopt a decoupled plug-in paradigm for semantic control over emotion and pose, enabling extensibility without disrupting the learned facial dynamics. Extensive experiments demonstrate that 3DXTalker achieves strong overall performance in identity preservation, lip synchronization, emotional expression, and controllable pose dynamics in 3D talking avatar generation.
Figure 1: Overview of the proposed expressive 3DXTalker system with plug-in semantic control. Given a static reference image and a driving audio, 3DXTalker generates expressive 3D talking avatars with consistent identity, accurate lip synchronization, rich emotional expression, and controllable head pose dynamics. Demo videos are provided in the supplementary material.
@@ 1 Introduction
Audio-driven 3D talking avatars have been widely applied across various domains [61, 4, 55, 28] by mapping audio signals (speech or song) to realistic 3D facial movements, providing an effective way to animate avatars without complex 3D capture hardware. Early approaches [16, 12, 41] focused on producing lip movements, which underperformed against the emerging demand for personalized and expressive avatars. As applications evolve, this field has shifted toward a more comprehensive goal: achieving greater Expressivity, where avatars could preserve identity, synchronize lip motion with speech, convey emotional expressions, and exhibit pose dynamics, while supporting global semantic control, as illustrated in Figure 1.
Despite notable progress [14, 37, 27, 45], achieving the goal of comprehensive expressivity in 3D talking avatars remains an ongoing challenge due to: 1) insufficient dataset coverage of diverse facial attributes; 2) the absence of an integrated framework for joint modeling of identity consistency, emotional expression, and pose dynamics; and 3) limited semantic controllability for flexible global style modulation. Specifically, 3D audio–mesh datasets [12, 17, 53, 37, 51] rely on costly real-world motion-capture devices, while current 2D-to-3D reconstructed datasets [45, 52] lack diverse identities and facial detail, hindering generalization. This limited coverage also impedes the creation of an integrated framework, while high-quality 3D avatars require a seamless fusion of expressivity from reference images and driving audio. Moreover, real-world avatar systems require broader semantic controllability for style modulation (e.g., emotion and pose control), but the open-ended nature of such demands makes continuous data collection and retraining impractical. These limitations hinder a model’s ability to capture the full spectrum of speech-driven expressivity for 3D avatars.
Following 2D-to-3D lifting paradigms [45, 52], we seek to address insufficient dataset coverage by constructing a scalable bridge from large-scale 2D videos to structured 3D facial representations, analogous to the VAE in text-to-image generation. Built upon FLAME [29] and DECA [18], EMOCA [13] recovers disentangled shape, expression, and pose parameters, along with fine facial details from video frames, making it suitable for our 3D facial modeling (Details in Part 1 of Sec. 2.). Thus, we construct a 2D-to-3D data integration pipeline over three lab-controlled datasets and three in-the-wild datasets, with statistics in Figure 2 (b) and (c). After unified filtering on duration, language, synchronization, and resolution, we build structured identity-aware FLAME representations via EMOCA and separate identity attributes from facial motions. This yields a dataset with rich identity coverage, lip and emotional variations, flexible pose dynamics, without requiring expensive 4D scanning.
@@
Utilizing the above dataset, we further introduce 3DXTalker, an integrated flow-matching framework that operates in the disentangled parameter space to generate expressive 3D talking-head motion sequences conditioned on a reference image and driving audio. Beyond conventional audio embeddings for lip sync, we incorporate frame-wise amplitude features for coherent mouth aperture and frame-wise emotion features for nuanced expression modulation. This forms audio-rich representations that more faithfully reflect the dynamics of speech. The image-derived identity latent and audio-derived motion cues are jointly modeled via a multi-branch transformer, which produces identity-consistent, emotionally aligned talking-head motion. To further enhance semantic controllability, we leverage the disentangled flow-matching framework to adopt a decoupled plug-in paradigm, which is instantiated with two representative examples, emotion control and head-pose control at inference time, to demonstrate its extensibility while preserving learned facial dynamics. With these advances, 3DXTalker provides a unified paradigm for expressive 3D talking avatars, simultaneously addressing identity consistency, lip-sync accuracy, rich emotion, and flexible pose dynamics.
The main contributions of this paper are summarized as follows:
@@
    We construct an identity-aware 3D talking dataset that bridges large-scale videos and structured 3D representations, improving attribute diversity to support more diverse avatars.
@@
    We introduce a flow-matching framework with audio-rich representations that incorporate frame-wise amplitude and emotion features, improving holistic expression fidelity.
@@
    We propose a decoupled plug-in paradigm for inference-time semantic control, which avoids retraining for open-ended demands and provides extensible controllability for 3D talking avatars.
Figure 2: Overview of the proposed 3DXTalker dataset. (a) Comparison of three typical 3D talking-head representation paradigms. (b) Diverse sources of the 3DXTalker dataset from both lab and wild environments, together with a rich word-cloud text corpus. (c) Comparison of existing datasets in subject and sample scale, and attribute coverage, highlighting the suitability of the 3DXTalker dataset for expressive 3D talking avatar. Details of datasets are presented in Appendix B.1.
@@ 2 Related Work
3D Facial Modeling and Datasets. The goal of audio-driven 3D avatar synthesis is to generate realistic and expressive 3D facial motion from speech. While early methods mainly relied on expensive motion-capture setups [12, 17, 53, 37, 51], recent advances in monocular 2D-to-3D face reconstruction have enabled direct lifting from images and videos to structured 3D facial feature representations. Existing methods mainly differ in their 3D facial representations, which can be broadly grouped into parametric and non-parametric forms. Parametric models, such as 3DMM [5], BFM [21], and FLAME [29], represent facial geometry with low-dimensional shape, expression, and pose parameters, making them widely used in 3D facial modeling [18, 13, 20, 59]. In particular, FLAME-based methods [40, 60, 19, 23] can recover disentangled shape, expression, and pose parameters, providing a structured representation for 3D facial animation. Building on this, EMOCA [13] further models fine facial details beyond standard FLAME parameters, enabling more faithful identity preservation, which makes it especially suitable for expressive 3D talking avatar generation. Non-parametric approaches [35, 58, 50, 8] model 3D facial mesh animations by directly learning per-vertex displacement [12, 41] or blendshape linear weighting [37, 15]. However, as shown in Figure 2 (a), such representations often remain high-dimensional or semantically entangled, which makes fine-grained motion control more difficult despite their strong deformation flexibility. More importantly, existing 2D-to-3D reconstructed datasets [45, 52] still lack sufficient comprehensive coverage of both identity-related geometry and fine facial details for expressive 3D talking avatars. We therefore adopt EMOCA as our 2D-to-3D bridge, playing a role similar to a VAE in text-to-image generation by encoding input facial frames into a compact yet expressive latent representation that supports expressive avatar generation.
@@
Audio-driven Regression Models. Early methods mainly formulate audio-driven 3D facial animation as a regression problem, where speech features are deterministically mapped to facial motion. These approaches typically employ features extracted from large-scale self-supervised speech models, such as Wav2vec 2.0 [3], HuBERT [24], and WavLM [7], and regress either raw vertex displacements [16, 34, 36, 26] or parametric model latents [27, 33, 46, 15, 54, 37, 14, 42]. FaceFormer [16] predicts vertex trajectories using a transformer decoder, while UniTalker [15] and CodeTalker [54] compress facial motion into low-dimensional latents for efficient learning. Although regression models can achieve good lip synchronization, their deterministic formulation is less effective at modeling the inherently one-to-many nature of speech-driven facial motion, often leading to limited expressivity and reduced facial dynamics that generative modeling can better support.
@@
Audio-driven Generative Models. To improve the diversity and realism of audio-driven 3D facial animation, recent works increasingly adopt generative formulations that model the conditional distribution from speech audio to facial motion. In particular, diffusion-based approaches (e.g., FaceDiffuser [44], DiffPoseTalk [45], DiffusionTalker [6], FaceTalk [1]) generate motion by progressively denoising from Gaussian noise, yielding more diversity and realism. Among them, some works mainly focus on lip synchronization, while others emphasize emotional facial expressions or head-pose dynamics separately. For instance, EMOTE [14] focuses on modeling emotional facial expressions, while DiffPoseTalk [45] explores audio-driven head-pose dynamics as a separate component. Despite these improvements in expressivity, existing methods still lack an integrated framework that jointly addresses identity consistency, expression, pose, and detail, especially for unseen subjects. Therefore, we present an integrated flow-matching framework with audio-rich representations and decoupled plug-in semantic control for expressive 3D talking avatar generation.
@@ 3 Method
In the task of audio-driven 3D talking avatar generation, the core objective is to synthesize a sequence of 3D avatar states $\{\mathbf{M}_{i}\}_{i=1}^{N}$ (NN total frames) that align with a given audio waveform $\mathbf{A}$, ensuring precise synchronization between facial movements and speech. Beyond lip synchronization, expressive avatar generation additionally requires identity preservation, emotion-aware facial expressions, and natural pose dynamics. However, this remains challenging due to limited training data with diverse facial attributes, the lack of an integrated framework for expressive facial dynamics, and restricted semantic controllability. To address these issues, we propose 3DXTalker, which improves expressive avatar generation through data integration and reformulation, the integrated flow-matching framework, and decoupled plug-in semantic control as follows.
@@
3.1 Data Integration and Reformulation
Scalable Data Integration. To overcome the limited attribute coverage of existing 3D talking-avatar datasets while avoiding expensive 3D capture, we built a scalable data integration pipeline that unifies diverse video sources into a structured 3D facial representation space. As illustrated in Figure 2 (b) and (c), we integrate three lab-controlled datasets (GRID [11], RAVDESS [31], MEAD [49]) and three in-the-wild datasets (VoxCeleb2 [10], HDTF [56], CelebV-HQ [57]) into a shared 3D representation space. The lab-controlled datasets provide clean articulation and explicit emotional variation, while the in-the-wild datasets contribute broader identity diversity, speaking styles, and natural pose dynamics.
To improve reliability and consistency across heterogeneous sources, we apply a unified filtering pipeline that consists of: (1) duration thresholding to retain or stitch clips longer than 10 seconds, (2) signal-to-noise ratio filtering to suppress noisy or corrupted speech, (3) language filtering to maintain linguistic consistency, (4) audio–visual synchronization verification to remove misaligned segments, and (5) spatial resolution normalization to $512\times 512$. Through this process, large-scale 2D videos are converted into a structured corpus suitable for expressive 3D talking-avatar modeling. Further details are provided in Appendix B.1.
Identity-Aware Differential Reformulation. After data integration, each video frame is lifted into the FLAME parameter space using the EMOCA encoder [13], yielding a structured representation, including shape $\boldsymbol{\beta}\in\mathbb{R}^{100}$, pose $\boldsymbol{\theta}\in\mathbb{R}^{6}$, expression $\boldsymbol{\psi}\in\mathbb{R}^{50}$ and an additional detail parameter $\boldsymbol{\delta}\in\mathbb{R}^{128}$. Notably, the pose $\boldsymbol{\theta}$ can be decomposed into head pose and jaw pose. These parameters can be decoded into a coarse mesh $\mathbf{M}_{coa}\in\mathbb{R}^{5023\times 3}$ deformed from a FLAME template $\mathbf{\bar{V}}$ with $\boldsymbol{\beta}$, $\boldsymbol{\psi}$, and $\boldsymbol{\theta}$, which is refined by the detail decoder $\mathcal{D}_{det}$} to form a detailed mesh $\mathbf{M}_{det}\in\mathbb{R}^{59315\times 3}$ through a facial displacement map, formally defined as:
@@ 3.1 Data Integration and Reformulation
where $\mathbf{V}_{int}$ is an intermediate mesh, $\text{B}_{S}$ and $\text{B}_{E}$ are shaping and expressing functions, LBS denotes the linear blend skinning algorithm with joint regressors $\text{J}_{P}$ and skinning weights $\mathcal{W}$, $\mathbf{\bar{U}}$ denotes template UV map, and $\text{F}_{det}$ applies the predicted displacement map to the coarse mesh $\mathbf{M}_{coa}$. Details are provided in Appendix A.
Rather than modeling these frame-wise parameters directly, we further reformulate them into an identity-anchored differential representation to separate stable identity attributes from temporally varying facial motion. Concretely, we take the first frame as a reference and express subsequent frames in relative form:
@@
3.2 Integrated Flow-matching Framework
The data integration and reformulation pipeline obtains disentangled identity and motion representations from 2D videos using EMOCA. This allows a single reference image to serve as a reliable identity anchor during generation, in contrast to prior 3D audio-driven methods that relied on a fixed identity template. Additionally, the audio side remains a limiting factor. Speech signals inherently contain multiple layers of information: (1) linguistic content (word, phoneme), (2) articulatory dynamics reflected in amplitude and rhythm that drive jaw motion and mouth aperture, and (3) emotional prosody conveyed through intonation, energy contours, and vocal timbre. Common audio embeddings focus on linguistic content [3, 24, 7], but overlook prosodic cues, leading to correct word synchronization but weak mouth aperture and flat emotion expressions.
Integrated audio-rich Backbone. To address the aforementioned gap, we introduce an integrated flow-matching framework that operates in the disentangled parameter space to generate expressive 3D talking-head sequences conditioned on a reference image $\mathbf{I}_{0}$ and driving audio $\mathbf{A}$. Beyond conventional audio embeddings, we incorporate frame-wise amplitude features for coherent mouth aperture and frame-wise emotion features for nuanced expression modulation, forming audio-rich representations that more precisely reflect the dynamics of speech. The identity latent derived from the reference image and the audio cues is then jointly modeled through a multi-branch flow-matching transformer, enabling identity consistency, lip synchronization, and emotional alignment in talking-head motion, as depicted in Figure 3. We next provide the detailed process of this framework.
Given an input image $\mathbf{I}_{0}$, we first extract the reference parametric representation $\mathbf{X}_{ref}=(\boldsymbol{\beta}_{0},\boldsymbol{\theta}_{0},\boldsymbol{\psi}_{0},\boldsymbol{\delta}_{0})\in\mathbb{R}^{1\times 284}$, which is then combined with $t$ step-dependent noise $\boldsymbol{\varepsilon}_{t}\in\mathbb{R}^{N\times 284}$ (initial random noise $\boldsymbol{\varepsilon}_{0}$) to form the model’s input state $\mathbf{\tilde{X}}_{t}$ after the MLP layers. The input audio $\mathbf{A}$ is processed into linguistic embeddings $\mathbf{A}_{feat}\in\mathbb{R}^{N\times d}$ ($d$ represents the dimension) using WavLM [7]. In the Transformer backbone $\mathcal{D}_{flow}$, $\mathbf{\tilde{X}}_{t}$ serves as queries, while the audio embeddings $\mathbf{A}_{feat}$ act as keys and values. Through self-attention and cross-attention, the model produces the intermediate latent representation $\mathbf{H}_{t}$. The overall process can be summarized: